A box of material X changes volume from 8.75 m³ to 10.0 m³ , when temperature is raised from 42.0 °C to 144 °C. Mass of the box is 12.0 kg.


(a) Calculate the volume expansion coefficient of the material.


(b) Calculate the change in density of the box, if temperature changes by 248 °C and initial volume is 10.0 m³.

1
Expert's answer
2022-01-16T13:11:35-0500

"(a)\space\gamma=\dfrac{V_f-V_i}{V_i\Delta T}=1.4·10^{-3}/\text K.\\\space\\
(b)\space \Delta \rho=\dfrac m{V_f}-\dfrac m{V_i}=-0.17\text{ kg/m}^3."